OpenStudy (anonymous):

Find the partial derivative indicated. Assume the variables are restricted to a domain on which the function is defined. f(x,y)=(x^3)(e^sqrt(3xy)) Find fx

OpenStudy (anonymous):

On my attempt of this problem I got 3x^2e^(sqrt(3xy))+(x^3/(2sqrt(3xy))(e^sqrt(3xy)))

OpenStudy (zarkon):

\[f_{x}=\left[ 3x ^{2}+\frac{3x ^{3}y}{2\sqrt{3xy}} \right]e ^{\sqrt{3xy}}\]

OpenStudy (anonymous):

Did you use the product rule?

OpenStudy (zarkon):

yes...and chain rule

OpenStudy (anonymous):

...... so confusing

OpenStudy (anonymous):

That's the same that I had earlier, just as Zarkon.

OpenStudy (zarkon):

you were missing a 2

OpenStudy (anonymous):

Thanks for helping